OpenStudy (anonymous):

what is 3x/x+2<=1/x+2

OpenStudy (dumbsearch2):

I <3 the shire :3

OpenStudy (dumbsearch2):

haha :) And did you enter your question wrong? There is no solution, I can only simplify it.

OpenStudy (anonymous):

\[\frac{3x}{x+2}<\frac{1}{x+2}\iff \frac{3x-1}{x+2}<0\] is a start

OpenStudy (anonymous):

<=